RC Meeting and Run Plan, Thursday, December 5th 2019

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links


  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in or comes out of laser-controlled access.
  • Make sure beam-size is between 200-300um
  • Still many TO day/swing shifts available! You don't need TO training, and will be guided on-site by the target experts!
  • Raster: Will need to mask MIN raster current in FSD when running 1-2 pass, ask PD if we need permission/signature from someone.
    • UPDATE: William set min current alarm value to 0, effectively disabling the alarm. Unfortunate side effect is that we won't have an FSD for raster failure anymore.

Hall C

Ongoing Issues

  • Major issue with magnet GUI, Steve Lassiter is fixing with Rockwell [https://logbooks.jlab.org/entry/3747133
  • HMS Dipole NMR gui screwed up, no way to set target B [1]
    • Checked today and NMR was disconnected -> fix ongoing?
  • Helium leak in Moller warm return, (Dave Gaskell) [2]
    • Moller not finished warming up, expected we can start repairs tomorrow morning. [3]
    • Moller commissioning after repairs and cool-down (expected Monday)
  • Operational restrictions on optics target inconsistent with empty target (Brad/Greg)
    • Need to decide raster requirements
  • Raster (Mark)
    • needs to be setup for 2-pass (Bill Gunning, ongoing)
    • Anything else?
  • Timing cuts and pre-run calibration (Eric, ongoing)

Resolved Issues

  • Alarm handler properly connecting EPICS again (Steve Wood) [4].
  • SHMS Shower gain match and HV good-to-go [5]
  • Ion chamber calibration procedure determined (Lester, Tom, Brad, Dave Mack)
  • Minor alarms on low LH2
  • Lost EPICS readback on magnets [6]
    • Mike Fowler performing maintenance, so not an issue [7]


Ongoing Issues

  • Transverse laser lines:
    • Alignment done, polarization measured [8]
    • Power test will be done using same lasers as longitudinal test tonight
    • Other 4 lasers can be tested and commissioned without Hall Access
  • RTDs
    • Fabrication complete, to be installed tonight (and readout tested)
    • Major temperature fluctuations measured [9]
  • Pickup coil needs to be moved away from ladder today [10]
  • Target heating system
    • operational, but slightly under-performing [11]
  • Plethora of software and readback issues, can be addressed without hall access [12]
  • Safety check on light-tightness of enclosure
    • What procedure will we follow?

Resolved Issues

  • Target alignment withing spec, no further work needed (big milestone!) [13]
  • Longitudinal laser lines tested to full power, no issues [14]

Run Plan

Scenario 1: ready Friday evening

assuming target work finished by Friday

  • 2-pass commissioning over the weekend (through Monday)
  • If Moller ready by Tuesday: participate in spin-dance at 5-pass
  • Then go to 1-pass
  • TODO: precise settings for calorimeter checkout

Scenario 2: ready after weekend

assuming we encounter critical issues that prevent us to close the hall

  • Participate in spin-dance at 5-pass if Moller ready
  • go directly to 1-pass for all commissioning
  • TODO: need to share contingency all-1-pass run-plan, especially because we will be hard-pressed to get more than 3 days...

Plan through Friday

  • Hall work during Thursday day
  • Target work during night Thursday-Friday
  • Finish up hall work during Friday day, and engage interlocks on target.
  • Repair Moller during Friday day.
  • Get required signatures on ERR to allow for beam delivery on Friday (crucial software needs to be done!)
  •  ?: Ask day shift workers to come in around noon?
  • Need to move spectrometers with Steve Lassiter to go to 30degrees for scenario 1
  • Bullseye scan plans? [15]
  • Tentative run checklist for 2-pass startup [16]


Special Requests

  • Please be gentle with our target :-)